The members of the science club agreed to contribute $480. if there had been 3 more members in the club each member would have had to pay $8 dollars less. how many members were in the science club originally.
Answer by lwsshak3(11628)

The members of the science club agreed to contribute $480. if there had been 3 more members in the club each member would have had to pay $8 dollars less. how many members were in the science club originally
*
let x=number of members in the science club originally
480/x=original club dues
480/(x+3)=club dues with 3 more members
..
480/x-480/(x+3)=8
LCD:x(x+3)
480(x+3)-480x=8x(x+3)
480x+1440-480x=8x^2+24x
8x^2+24x-1440=0
x^2+3x-180=0
factor
(x+15)(x-12)=0
x=-15 (reject, x>0)
x=12
x+3=15
number of members in the science club originally=12